OpenStudy (blockcolder):

A little question on cross product. For vectors \(\mathbf{a},\mathbf{b} \in V_3\), does the property go \(\mathbf{a} \times \mathbf{b}=-\mathbf{b} \times \mathbf{a}\) or \(\mathbf{a} \times \mathbf{b}=-(\mathbf{b} \times \mathbf{a})\)?

mathslover (mathslover):

a*b = -ab or ab = -(ba) .... this is false .

OpenStudy (unklerhaukus):

the two if the question are equivalent

OpenStudy (anonymous):

same thing u wrote... nothing change

OpenStudy (unklerhaukus):

\[\alpha \mathbf a \times \mathbf b=\alpha( \mathbf a\times \mathbf b)\] where \(\alpha\in\mathbb R\)

OpenStudy (blockcolder):

Ah, so in this case, \(\alpha=-1\). Thanks for that! I was a little confused.

OpenStudy (unklerhaukus):

that is right

OpenStudy (anonymous):

agree wid unkle, the "-"sign will shift with wid.... its hvnt any problm.... all same

OpenStudy (anonymous):

with "b"... sorry*

OpenStudy (maheshmeghwal9):

http://www.physicsforums.com/showthread.php?t=318030